# Investigating FRB 20240114A with FAST: Morphological Classification and Drifting Rate Measurements in a Burst-Cluster Framework

### v1

@@ 4.3 Analysis of consecutive and intermittent for downward and upward drifting
Figure 13: Overall view of the sub-burst interval distributions in four panels. The top row displays the histogram and cumulative distribution for downward sub-burst intervals, whereas the bottom row shows the histogram and cumulative distribution for upward sub-burst intervals. These four panels together illustrate the statistical properties of sub-burst intervals for both drifting directions. Empirical CDFs of downward and upward sub-burst intervals with the K-S distance indicated.
@@
In this section, we present the statistical distribution of consecutive and intermittent time intervals for upward and downward drifting burst-clusters, derived from double-burst-cluster and multi-burst-cluster events. First, we focus on the intermittent intervals for upward and downward drifting burst-clusters separately. For the downward drifting burst-clusters, the sample size is 265, with a mean of 64.48 ms and a standard deviation of 89.40 ms, while for upward drifting burst-clusters, the sample size is 90, with a mean of 93.55 ms and a standard deviation of 124.91 ms. A Kolmogorov-Smirnov (K-S) test was performed on the intermittent only data, resulting in a test statistic D = 0.132 and a p-value of $1.740\times 10^{-1}$, as shown in the first row of Figure 13, which suggests that there is no statistically significant difference between the two types of drifting burst-clusters in terms of their intermittent intervals.
@@
Due to the scarcity of consecutive time intervals for upward drifting burst-clusters, with only 9 such intervals observed (shown in Figure 6), we were unable to perform a separate statistical analysis for consecutive upward and downward drifting burst-clusters. As a result, we combined consecutive and intermittent intervals for a comprehensive analysis. For the downward drifting burst-clusters, the combined sample size is 449, with a mean of 39.45 ms and a standard deviation of 74.94 ms, and for upward drifting burst-clusters, the combined sample size is 99, with a mean of 85.36 ms and a standard deviation of 121.85 ms. The K-S test for the combined data, shown in the second row of Figure 13, yielded a test statistic D = 0.314 and a p-value of $1.376\times 10^{-7}$, indicating that the upward and downward drifting burst-clusters come from different distributions.
@@
The analyses above indicate that there are statistically significant differences in the consecutive time intervals between upward and downward drifting burst-clusters. Since the intermittent upward and downward drifting burst-clusters come from the same distribution and significant differences were observed in the combined data of both intermittent and consecutive intervals, it suggests that the mechanisms driving consecutive upward and downward drifting burst-clusters are different. Specifically, upward drifting burst-clusters exhibit a much longer mean interval, which suggests distinct physical processes or factors governing these burst-clusters compared to downward drifting burst-clusters. Further investigation is necessary to better understand the nature of these processes and their impact on drifting behavior.
5 Conclusion and Discussion
@@
This study presents morphological classification and drifting rate measurements of the repeating fast radio burst source FRB 20240114A, utilizing observational data obtained from the Five-hundred-meter Aperture Spherical Telescope (FAST) on March 12, 2024. As detailed in Section 2.2, we employed the average dispersion measure (DM) optimized for the burst’s finest temporal structures during the March 12, 2024 observations to de-disperse the data. This enabled comprehensive morphological classification (Section 3) and drifting rate measurements (Section 4.1), followed by systematic exploration of statistical relationships between drifting rates and various physical parameters (Section 4.2). Key concepts including sub-bursts, bursts, burst-clusters, and consecutive/intermittent time intervals were rigorously defined and visually illustrated (Figure 3), with consecutive time intervals of downward/upward drifting burst-clusters (i.e., time intervals of downward/upward drifting bursts) also being specified. Notably, as discussed in Section 4.3, a Kolmogorov-Smirnov (K-S) test was implemented to statistically compare whether the intermittent and consecutive time intervals of upward/downward drifting burst-clusters originate from identical distributions.
In the analysis of the upward drifting single-component burst-clusters (U1), we observe a linear increase in the lower bound of the drifting rate as the peak frequency increases (Figure 10). This trend may be influenced by the accuracy of the DM correction. If the DM were perfectly corrected for a single-component burst-cluster, leading to ND1, inaccuracies in the DM correction could cause a higher drifting rate for high-frequency burst-clusters compared to low-frequency ones, making the drifting less obvious. However, no similar pattern is observed for downward drifting, suggesting that DM correction alone cannot fully explain the observed drifting behavior. This implies that other physical mechanisms, in addition to DM correction, may contribute to the drifting rate behavior.
@@
In our study of FRB 20240114A, we detect a small fraction of upward drifting burst-clusters (23.82%), decreasing to 10.89% when single upward drifting burst-clusters are excluded. Moreover, when only upward drifting burst-clusters consisting solely of consecutive time intervals are considered, this number further drops to just 9, as illustrated in Figure 6. Notably, these 9 consecutive upward drifting burst-clusters represent a proportion nearly identical to that reported for FRB 20201124A in Zhou et al. (2022), suggesting similar underlying physical mechanisms across different repeating FRB sources.
@@
The analysis presented in this work reveals a robust inverse correlation between the drifting rate ($R_{d}$) and the sub-burst effective widths (see Figure 10). This inverse relationship corroborates the findings of Chamma et al. (2021) and suggests that the dynamical processes governing the drifting phenomena are closely linked to the temporal characteristics of the burst emission. In addition, for single-component burst-clusters, our results show that downward drifting events exhibit larger bandwidths, effective widths, and flux compared to upward drifting events (see Figures 10 and 11). However, the physical origin underlying these observational differences remains unclear, and further investigations are necessary to fully understand the mechanisms responsible for these drifting behaviors.
@@
A Kolmogorov-Smirnov (K-S) test reveals that the consecutive time intervals of upward drifting burst-clusters are significantly longer than those of downward drifting burst-clusters (Figure 13), whereas the intermittent time intervals of upward/downward drifting burst-clusters are drawn from the same distribution, indicating that while intermittent time intervals of upward/downward drifting burst-clusters may result from random triggering, consecutive drifting behaviors reflect distinct physical mechanisms. Consequently, we concentrate on the drifting patterns of individual bursts. Fully understanding these mechanisms will require high-time-resolution follow-up observations and the development of more comprehensive theoretical models to explain the frequency drifting phenomena observed in FRB 20240114A.
@@
Most bursts exhibit downward drifting patterns, consistent with many repeating FRBs previously studied (e.g. Pleunis et al., 2021; Zhou et al., 2022). These downward drifting bursts are often interpreted within models invoking radius-to-frequency mapping in magnetospheric emission regions (Cordes & Chatterjee, 2019; Wang et al., 2019) or plasma lensing effects (Cordes et al., 2017; Tuntsov et al., 2021), where higher frequencies arrive prior to lower frequencies. This interpretation is supported by the even shorter and more numerous consecutive time intervals of downward drifting burst-clusters, as well as the well-defined negative drifting rates observed. Nevertheless, the morphological asymmetries between upward and downward drifting bursts could arise from the geometric effect of ordered bunch emission along different magnetic field lines. As proposed by Wang et al. (2022), the drifting direction (upward or downward) is determined by the sequence in which bunches cross the line of sight, with the curvature of field lines and the observer’s viewing angle playing crucial roles in shaping the distinct spectro-temporal patterns. The observed downward drifting patterns also find theoretical grounding in magnetospheric curvature radiation models, where relativistic particle bunches propagating along neutron star open field lines naturally exhibit downward drifting due to increasing curvature radii at higher altitudes (Wang et al., 2019). Notably, Wang et al. (2020) extended this framework to incorporate upward drifting scenarios, demonstrating that earlier-triggered sparks at larger emission heights can reverse the drifting direction under specific geometric configurations and temporal sequencing of magnetospheric perturbations. Our detection of upward drifting burst-clusters (23.82% overall, with only nine exceptional consecutive cases) aligns with this extended framework, suggesting rare instances where spark triggering sequences or non-dipolar field geometries dominate over the standard downward drifting mechanism.
@@ Acknowledgments
This work made use of the data from FAST FRB Key Science Project. This work is supported by National Natural Science Foundation of China (NSFC) Programs Nos. 12588202, 12041306, 11988101, 11725313, 11690024, 12041303, 12421003, U1731238, 12233002, 12041306, 12303042, 12403100, W2442001, 12203045, 12447115 and U1931203; CAS International Partnership Program No. 114-A11KYSB20160008; CAS Strategic Priority Research Program No. XDB23000000; the National Key R&D Program of China (No. 2017YFA0402600); the National SKA Program of China Nos. 2020SKA0120200, 2020SKA0120100 and 2022SKA0130100. W.W.Z. is supported by CAS Project for Young Scientists in Basic Research, YSBR-063. D.L. is a New Cornerstone investigator. P.W. acknowledges support from the CAS Youth Interdisciplinary Team, the Youth Innovation Promotion Association CAS (id. 2021055), and the Cultivation Project for FAST Scientific Payoff and Research Achievement of CAMS-CAS. Y.F. is supported by the Leading Innovation and Entrepreneurship Team of Zhejiang Province of China grant No. 2023R01008, and by Key R&D Program of Zhejiang grant No. 2024SSYS0012. Y.F.H. acknowledges the support from the Xinjiang Tianchi Program. Q.W. is supported by the China Postdoctoral Science Foundation (CPSF) under Grant Nos. GZB20240308, 2025T180875, 2025T180875. C.W.T. is supported by CAS project No. JZHKYPT-2021-06. This work made use of data from FAST, a Chinese national mega-science facility built and operated by the National Astronomical Observatories, Chinese Academy of Sciences. This research has made use of the CHIME/FRB VOEvent Service, BlinkVerse and TransientVerse. The cartoon plot was created with the help of Ms. Yan-Dong Wang and Mr. Tong-Cai Zhang. The computation was partially completed on the HPC Platform of Huazhong University of Science and Technology.

### v3

@@ IV.3 Analysis of consecutive and intermittent for downward and upward drifting
Figure 14: Overall view of the sub-burst interval distributions in left four panels. Empirical CDFs of downward and upward sub-burst intervals with the K-S distance indicated in the right two panels. The top row displays the histogram and cumulative distribution for intermittent sub-burst intervals, whereas the bottom row shows the histogram and cumulative distribution for both intermittent and consecutive sub-burst intervals. These left four panels together illustrate the statistical properties of sub-burst intervals for both downward and upward drifting directions.
@@
In this section, we present the statistical distribution of consecutive and intermittent time intervals for upward and downward drifting burst-clusters, derived from double- and multi-component burst-clusters. First, we focus on the intermittent intervals for upward and downward drifting burst-clusters separately. For the downward drifting burst-clusters, the sample size is 265, with a mean of 64.48 ms and a standard deviation of 89.40 ms, while for upward drifting burst-clusters, the sample size is 90, with a mean of 93.55 ms and a standard deviation of 124.91 ms. A Kolmogorov-Smirnov (K-S) test was performed on the intermittent only data, resulting in a test statistic D = 0.132 and a p-value of $1.740\times 10^{-1}$, as shown in the first row of Figure 14. This high p-value indicates that we cannot reject the null hypothesis, meaning the intermittent intervals for the two types of drifting burst-clusters are statistically consistent with being drawn from the same distribution.
@@
Due to the scarcity of consecutive time intervals for upward drifting burst-clusters, with only 9 such intervals observed (shown in Figure 6), we were unable to perform a separate statistical analysis for consecutive upward and downward drifting burst-clusters. As a result, we combined consecutive and intermittent intervals for a comprehensive analysis. For the downward drifting burst-clusters, the combined sample size is 449, with a mean of 39.45 ms and a standard deviation of 74.94 ms, and for upward drifting burst-clusters, the combined sample size is 99, with a mean of 85.36 ms and a standard deviation of 121.85 ms. The K-S test for the combined data, shown in the second row of Figure 14, yielded a test statistic D = 0.314 and a p-value of $1.376\times 10^{-7}$. This very small p-value provides strong evidence to reject the null hypothesis, suggesting that the combined intervals for upward and downward drifting burst-clusters are likely drawn from different distributions.
@@
The analyses above indicate a statistically significant difference ($p<0.05$) in the combined (consecutive and intermittent) time intervals between upward and downward drifting burst-clusters. Since the intermittent upward and downward drifting burst-clusters come from the same distribution and significant differences were observed in the combined data of both intermittent and consecutive intervals, it suggests that the mechanisms driving consecutive upward and downward drifting burst-clusters are different. Specifically, upward drifting burst-clusters exhibit a much longer mean interval, which suggests distinct physical processes or factors governing these burst-clusters compared to downward drifting burst-clusters. Further investigation is necessary to better understand the nature of these processes and their impact on drifting behavior.
V Conclusion and Discussion
@@
This study presents the morphological classification and drifting rate measurements of the repeating fast radio burst source FRB 20240114A, utilizing observational data obtained from FAST on March 12, 2024. As detailed in Section II.2, we employed the average DM optimized for the burst’s finest temporal structures during the March 12, 2024, observations to de-disperse the data. This enabled comprehensive morphological classification (Section III) and drifting rate measurements (Section IV.1), followed by a systematic exploration of statistical relations between drifting rates and various physical parameters, as well as the narrow-band spectral properties that are more naturally explained by magnetospheric emission (Section IV.2). Key concepts, including sub-bursts, bursts, burst-clusters, and consecutive/intermittent time intervals, were rigorously defined and visually illustrated (Figure 3), with consecutive time intervals of downward/upward drifting burst-clusters (i.e., time intervals of downward/upward drifting bursts) also being specified. Notably, as discussed in Section IV.3, a K-S test was implemented to statistically compare whether the intermittent and consecutive time intervals of upward/downward drifting burst-clusters originate from identical distributions.
@@
In our study of FRB 20240114A, we detected a small fraction of upward drifting burst-clusters (23.82%), decreasing to 10.89% when single-component upward drifting burst-clusters are excluded. Moreover, when only upward drifting burst-clusters consisting solely of consecutive time intervals are considered, this number further drops to just 9, as illustrated in Figure 6. Notably, these 9 consecutive upward drifting burst-clusters represent a proportion nearly the same as that reported for FRB 20201124A in 2022RAA....22l4001Z. Considering that the mechanism of upward drifting is still unclear (though a few models have been proposed, e.g. 2020ApJ...899..109W), such rare upward series in multiple FRBs might hint at analogous conditions in their sources.
In the analysis of the single-component burst-clusters, we observe a positive correlation between the drifting rate and the peak frequency (left panel of Figure 12). This trend may be attributed to deviations between the adopted average DM and the true per-burst DM. The theoretical derivation shows that, under the assumption that DM deviations are the only source of the observed slope, the measured drifting rate should increase with frequency. However, both U1 and D1 burst-clusters exhibit different linear correlations. This indicates that DM deviations alone cannot fully account for the observed drifting behavior, implying that additional, possibly unknown processes may contribute to the observed frequency dependence.
@@
The analysis presented in this work reveals an inverse correlation between the drifting rate and the sub-burst effective widths (see Figure  10). This inverse relation corroborates the findings of 2021MNRAS.507..246C, which may indicate a triggered scenario involving a compact object (e.g. a pulsar or magnetar) and a separate FRB-emitting region along the line of sight. For single-component burst-clusters, our results show that downward drifting events exhibit larger bandwidths, effective widths, and flux compared to upward drifting events (see Figures 10 and 12). However, the physical origin underlying these observational differences remains unclear, and further investigations are necessary to fully understand the mechanisms responsible for these drifting behaviors. In addition, the fractional bandwidths remain narrowly confined, with FWHM-corrected values of $\Delta\nu/\nu_{0}\approx 0.22$-$0.27$ (see Figures 13), well below the $\Delta\nu/\nu_{0}\sim 0.58$ limit expected for emission outside the magnetosphere (2024ApJ...974..160K). This trend is more naturally accommodated by magnetospheric scenarios, though propagation effects may still contribute. Similar narrow-band behavior has also been reported for other repeaters (2023ApJ...955..142Z; 2025ApJ...988...41Z; 2024ApJ...977..177K).
@@
As shown in Section IV.3, the consecutive time intervals of upward drifting burst-clusters (mean value 93.35 ms) are significantly longer than those of downward drifting burst-clusters (mean value 64.48 ms). In contrast, the intermittent time intervals of upward/downward drifting burst-clusters are drawn from the same distribution (as indicated by the K-S test shown in Figure 14), suggesting that while the intermittent time intervals of upward/downward drifting burst-clusters may result from random triggering, the consecutive drifting behaviors reflect distinct physical mechanisms, although it is not clear what those mechanisms should be. Consequently, we concentrate on the drifting patterns of individual bursts. Fully understanding these mechanisms will require high-time-resolution follow-up observations and the development of more comprehensive theoretical models to explain the frequency drifting phenomena observed in FRB 20240114A.
@@
Most bursts exhibit downward drifting patterns, consistent with many repeating FRBs previously studied (e.g. 2021ApJ...923....1P; 2022RAA....22l4001Z). These downward drifting bursts are often interpreted within models invoking radius-to-frequency mapping in magnetospheric emission regions (2019ARA&A..57..417C; 2019ApJ...876L..15W) or plasma lensing effects (2017ApJ...842...35C; 2021arXiv210713549T), where higher frequencies arrive prior to lower frequencies. This interpretation is supported by the even shorter and more numerous consecutive time intervals of downward drifting burst-clusters, as well as the well-defined negative drifting rates observed. Nevertheless, the morphological asymmetries between upward and downward drifting bursts could arise from the geometric effect of ordered bunch emission along different magnetic field lines. As proposed by 2022ApJ...927..105W, the drifting direction (upward or downward) is determined by the sequence in which bunches cross the line of sight, with the curvature of field lines and the observer’s viewing angle playing crucial roles in shaping the distinct spectro-temporal patterns. The observed downward drifting patterns also find theoretical grounding in magnetospheric curvature radiation models, where relativistic particle bunches propagating along neutron star open field lines naturally exhibit downward drifting due to increasing curvature radii at higher altitudes (2019ApJ...876L..15W). Notably, 2020ApJ...899..109W extended this framework to incorporate upward drifting scenarios, demonstrating that earlier-triggered sparks at larger emission heights can reverse the drifting direction under specific geometric configurations and the temporal sequencing of magnetospheric perturbations. Our detection of upward drifting burst-clusters (23.82% overall, with only nine exceptional consecutive cases) aligns with this extended framework, suggesting rare instances where spark triggering sequences or non-dipolar field geometries dominate over the standard downward drifting mechanism.
Data Availability
The dataset underlying this work is publicly available at the Science Data Bank https://doi.org/10.57760/sciencedb.Fastro.00033. It includes dynamic spectra of FRB 20240114A in archival format (.ar files), widely used in radio astronomy, along with their corresponding visualization plots (.png). The dataset also provides summary tables in .xlsx format, covering morphological classifications, drifting rate measurements, and other relevant physical quantities. In addition, curated data subsets are included for the K–S test. All files are provided in open formats to support transparent and reproducible research.
@@ Acknowledgments
This work made use of the data from FAST FRB Key Science Project. This work is supported by National Natural Science Foundation of China (NSFC) Programs Nos. 12588202, 12041306, 11988101, 11725313, 11690024, 12041303, 12421003, U1731238, 12233002, 12303042, 12403100, W2442001, 12203045, 12447115 and U1931203; CAS International Partnership Program No. 114-A11KYSB20160008; CAS Strategic Priority Research Program No. XDB23000000; the National Key R&D Program of China (No. 2017YFA0402600); the National SKA Program of China Nos. 2022SKA0130100, 2020SKA0120200, and 2020SKA0120100. W.W.Z. is supported by CAS Project for Young Scientists in Basic Research, YSBR-063. D.L. is a New Cornerstone investigator. P.W. acknowledges support from the CAS Youth Interdisciplinary Team, the Youth Innovation Promotion Association CAS (id. 2021055), and the Cultivation Project for FAST Scientific Payoff and Research Achievement of CAMS-CAS. Y.F. is supported by the Leading Innovation and Entrepreneurship Team of Zhejiang Province of China grant No. 2023R01008, and by Key R&D Program of Zhejiang grant No. 2024SSYS0012. Y.F.H. acknowledges the support from the Xinjiang Tianchi Program. Q.W. is supported by the China Postdoctoral Science Foundation (CPSF) under Grant Nos. GZB20240308, 2025T180875. C.W.T. is supported by CAS project No. JZHKYPT-2021-06. J.R.N. is supported by the Postdoctoral Fellowship Program of CPSF under Grant No. GZB20250737. This work made use of data from FAST, a Chinese national mega-science facility built and operated by the National Astronomical Observatories, Chinese Academy of Sciences. This research has made use of the CHIME/FRB VOEvent Service, BlinkVerse and TransientVerse. The cartoon plot was created with the help of Ms. Yan-Dong Wang and Mr. Tong-Cai Zhang. The computation was partially completed on the HPC Platform of Huazhong University of Science and Technology.